# A Physics-Constrained Neural Differential Equation Framework for Data-Driven Snowpack Simulation

### v2

@@ 1 Introduction
[…]
@@ 2.4 Data
The training data were averaged (preserving start-of-window $z$ and SWE) over a consecutive $N$-day moving window, with $N$ as a hyperparameter. This enabled exploration of the tradeoff of spreading out discretized SNOTEL data ($z$ to the inch, $\mathrm{SWE}$ to 0.1 inch) for smoother regression learning against preserving extreme values critical for predictions. Days with unphysical values (zero $\mathrm{SWE}$ and nonzero $z$) or no snowpack were removed to eliminate uninformative zeros in the target space. Features were scaled by their standard deviations, and the target by its absolute maximum, with scaling constants fixed in $M$ to spare user preprocessing. This resulted in 58,484 usable sensor-days out of 105,636 for training, and 35,618 sensor-days for testing. The complete dataset and generating code are publicly available (see Data Availability).
@@ 2.6 Testing
In addition to depth, bulk density timeseries can be generated from $z$ and SWE data and model outputs according to Eq. 1 and compared. The data is discrete while the model outputs are continuous, so densities were compared at sites with collocated $\mathrm{SWE}$ and $z$ sensors (see Appendix B) only when both models and data yielded physical values ($10<\rho_{\mathrm{snow}}/\rho_{\mathrm{water}}<$ 1). Counts of “false non-snowpacks” (models show $z=0$ while $z>0$ in the data) and “false snowpacks” (models show $z>0$ while $z=0$ in the data) were recorded, along with instances of unphysical densities.
@@
Assessing physical consistency in the model directly per feature is nuanced due to strong correlations among the inputs, which confounds most interpretation methods like partial dependence, permutation, LIME, and tractable SHAP values (Molnar, 2022). To isolate feature effects on model output, we calculated first-order Accumulated Local Effects (ALE) plots (Apley and Zhu, 2020). This method shows the average change in model output accumulated over sequential bins of feature values. For feature $X$ at grid point $x_{i}$, data in the window $[x_{i}-\Delta,x_{i}+\Delta]$ are evaluated setting $X=x_{i}-\Delta$ and $X=x_{i}+\Delta$, storing the average of the differences $\Delta M_{x_{i}}=\overline{M(x_{i}+\Delta)-M(x_{i}-\Delta)}$. The final (centered) ALE value $\bar{\Delta M}$ at $x_{i}$ is the sum over all $\Delta M_{x_{k}}$ with $x_{k}\leq x_{i}$, minus the average of all uncentered ALE values. This isolates changes solely from feature variations, and bins are defined by quantiles to ensure equal data instances in each window. The shape and slope of the ALE curve are more pertinent for interpreting physicality than the offset, especially when the feature distribution is skewed. The range of $\Delta\bar{M}$ over the feature indirectly measures feature importance in influencing model output, as the ALE value can be interpreted as a departure from the average model prediction. Further information for interpreting ALE plots can be found in Molnar (2022).
@@ 3.2 Density Timeseries
For validation and testing sites with collocated $z$ and SWE measurements (see Appendix B), Table 1 compares timeseries statistics for derived bulk density, and Fig. 7 displays estimates from the same data as Fig. 6. RMSE values reflect relative error since density is normalized by $\rho_{\mathrm{water}}$. Both SN17 and $\tilde{M}$ lack SWE discretization, inflating errors for small snowpacks compared to their assimilated counterparts. Neural unphysical densities source from a lack of constraint enforcing $\mathrm{SWE}=0\implies z=0$ (in both $\tilde{M},M$) or $z>\mathrm{SWE}$ (in $M$), though different constraint choices could eliminate this. Snow17 is roughly 3×3$\times$3 × better at mitigating false snowpacks with lower RMSE (less extreme errors), but our models are roughly 3×3$\times$3 × better at reducing false snow-absence and exhibit improved individual density predictions, crucial for estimating other snowpack properties. Thus, the advantage between models depends on which features are prioritized. Otherwise, the two perform similarly, despite $M$ being designed for snow depth while the comparable parameterization within Snow17 is explicitly formulated for density. Among evaluated sites, the best and worst performances by $M$ were better than those of SN17O.
@@ 3.4 Physical Behavior of Model
Figure 10: $M$ outputs for two snow states, with other inputs held constant. In each case one threshold ($dz>-z$ or $dz/dt<0$ for $P_{\mathrm{snow}}=0$) is visible. “Shaded” or “masked” pixels (partially greyed out) indicate areas of the visible parameter space with ranges not represented in the data, either from being unphysical or unobserved.
@@ 3.6 Finer Resolution Predictions
Time units only appear as rates in $M$ via precipitation, the lower bound, and the output. By predicting rates $dz/dt$ instead of accumulated $dz$, $M$ can be evaluated at varying time-steps without retraining, by merely resetting the constraint function’s scaling of $1/\Delta t$ without altering the trained predictive weights (which were trained only once in section 3a, with $\Delta t$ = 1 day). This flexibility allows testing at resolutions up to data limits, which for Kühtai is a maximum resolution of 15-minute intervals. While sub-hourly time resolutions are rarer in larger earth models, they are used in land models and will become increasingly frequent as research advances in fine-scale land-atmosphere interactions (Schär et al., 2020; Ban et al., 2021). Similarly, while multiday resolutions are rare, data availability can necessitate their use. Benchmarking across scales above and below the daily resolution used for training offers insights into the model’s limits and potential applications. However, beyond a week, average input variables do not effectively capture critical input dynamics (e.g., monthly average $T_{\mathrm{air}}$ would not reliably indicate time above freezing), leading to suppressed variations that undermine meaningful outputs, and alternative model choices would be more appropriate. Results of repeating the timeseries generation with $M$ as described in Section 3f for the Kühtai data with different values of $\Delta t$ in the constraint but identical weights otherwise, spanning 15-minute to weekly resolutions, are shown in Fig. 12 and Table 2.
@@ Acknowledgements.
Then for a model output $p$ and any construction $f$ serving to threshold $p$, the bounds $\max(f,p)$ or $\min(f,p)$ can be explicitly implemented with a single depth-3 fixed-weight layer with no biases acting on input $[p,f]^{\top}$ with ReLU activation, followed by an accumulation with no activation:
[±$\begin{bmatrix}\pm 1&1&-1\end{bmatrix}\times\mathrm{ReLU}\left(\begin{bmatrix}\pm 1&\mp 1\\ 0&1\\ 0&-1\end{bmatrix}\times\begin{bmatrix}p\\ f\end{bmatrix}\right)\\ =\bm{\mathsf{A}}_{1\pm}^{\top}\>\mathrm{ReLU}\left(\bm{\mathsf{A}}_{2\pm}^{\top}\begin{bmatrix}p\\ f\end{bmatrix}\right),$ (9)
where $+$ indicates $\max(f,p)$ and -- indicates $\min(f,p)$, and the ReLU acts element-wise. Eq. 9 offers a simple formulation, but the symmetry of the max and min functions permits resonant structures—different weights yielding the same output. If the signs of $p$ or $f$ are unknown, both $+f$ and $-f$ (or $+p$ and $-p$) must be passed through the ReLU, along with $p-f$, to preserve all necessary information. However, if $f$ is always nonnegative (or nonpositive), the layer can be simplified from three layers to two, as passing $-f$ (or $+f$) becomes redundant and its ReLU always evaluates to zero. Similarly, if the threshold obeys $f\geq C$ (or $p\geq C$) for some constant $C$, a similar reduction is possible by including a bias term along with $A_{2\pm}$ and $A_{1\pm}$. The same reductions apply if $p$ also exhibits similar properties. Fig. 13a illustrates the generalized structure for one-sided threshold constraints on a predictive component discussed in section 22.2.
@@
so the threshold can be explicitly implemented with a sequence of two fixed-weight layers containing no biases acting on input $[p,f_{+},f_{-}]^{\top}$, followed by an accumulation with no activation:
$\begin{bmatrix}1&1&-1\end{bmatrix}\times\mathrm{ReLU}\left(\begin{bmatrix}-1&1&-1&-1&1\\ 0&0&0&1&0\\ 0&0&0&0&1\end{bmatrix}\right.\\ \left.\times\;\mathrm{ReLU}\left(\begin{bmatrix}-1&1&0\\ 0&1&0\\ 0&-1&0\\ 0&0&1\\ 0&0&-1\end{bmatrix}\times\begin{bmatrix}p\\ f_{+}\\ f_{-}\end{bmatrix}\right)\right)$ (12)
$=\bm{\mathsf{A}}_{1+}^{\top}\>\mathrm{ReLU}\left(\bm{\mathsf{A}}_{3}^{\top}\mathrm{ReLU}\left(\bm{\mathsf{A}}_{4}^{\top}\begin{bmatrix}p\\ f_{+}\\ f_{-}\end{bmatrix}\right)\right),$ (13)
@@
These constraint structures can adapt to any functional constraint $f$ of any input (including those independent of predictive inputs) or even the predictive component output $p$. This versatile framework exhibits minimal computational overhead without increasing runtime complexity. This thresholds an output rather than imposing an invariance or zeroing of a derivative, but this constraint class is relevant to many physical or nonnatural systems. This provides a simple means for emulating many systems when paired with universally approximating predictive models, or integrating data-driven parameterizations into larger models while enforcing constraints like physics or conservation laws. Absolute boundaries also enhance stability in time-stepping by keeping outputs realistic.
@@
  • Only days with complete cases (no missing values in all variables) were extracted, and sequential differences $\Delta z_{i}=z_{i+1}-z_{i}$, $\Delta\mathrm{SWE}_{i}=\mathrm{SWE}_{i+1}-\mathrm{SWE_{i}}$, $p_{i}=AP_{i+1}-AP_{i}$ were calculated, where $AP$ is accumulated precipitation. Only values where $\Delta t_{i}=t_{i+1}-t_{i}=1$ day were kept, and the target $\frac{dz}{dt}_{i}\approx\frac{\Delta z_{i}}{\Delta t_{i}}$ created, as well as analogous $\frac{d\mathrm{SWE}}{dt}_{i}$ and $P_{i}=p_{i}/\Delta t_{i}$. Days with $P_{i}<0$ due to resetting of the water year were changed to $P_{i}$ = 0.
@@
  • Air temperatures in the continental US were bounded below by $-40^{\circ}$ and air temperatures in Alaska were bounded below by $-50^{\circ}$ instead of the instrument limit of $-60^{\circ}$ to remove on average 1-2 individual suspect temperature spikes per site.

### v3

@@ 1 Introduction
Modeling the evolution of seasonal snow for regional or global climate applications offers a challenging problem of scales; it is the bulk properties of the snow (albedo, snow cover fraction, snow temperature, and snow water content) that are critical, yet microphysical and location-specific processes control these properties and must be taken into account. The most detailed models represent vertically-resolved snowpacks, including liquid percolation, phase changes, metamorphism effects, and other types of compaction; they are often calibrated and used on the site-level (e.g., De Michele et al. 2013b). Models used in global climate simulations range in complexity from single-layer/bulk models to multi-layer models with parameterizations for bulk properties that are calibrated with observational data (Menard et al., 2021); the horizontal resolution of these models is typically ${\sim}10$100100 km where microscopic processes cannot be tractably resolved. While the laws of physics ultimately govern the evolution of snowpacks, uncertainty in how to relate essential but unresolved small-scale processes to snowpack bulk properties on the spatial scales of global models makes developing snow models a challenging task (Kapnick et al., 2018; Bair et al., 2018). This challenge is exacerbated by data availability (Menard et al., 2021; Kouki et al., 2022).
@@ 2.4 Data
The training data were averaged (preserving start-of-window $z$ and SWE) over a consecutive NN-day moving window, with NN as a hyperparameter. This enabled exploration of the tradeoff of spreading out discretized SNOTEL data ($z$ to the inch, $\mathrm{SWE}$ to 0.1 inch) for smoother regression learning against preserving extreme values critical for predictions. Days with unphysical values (zero $\mathrm{SWE}$ and nonzero $z$) or no snowpack were removed to eliminate uninformative zeros in the target space. Features were scaled by their standard deviations, and the target by its absolute maximum, with scaling constants fixed in $M$ to spare user preprocessing. This resulted in 58,484 usable sensor-days out of 105,636 for training, and 35,618 sensor-days for testing. The complete dataset and generating code are publicly available (see Data Availability).
@@ 2.6 Testing
In addition to depth, bulk density timeseries can be generated from $z$ and SWE data and model outputs according to Eq. 1 and compared. The data is discrete while the model outputs are continuous, so densities were compared at sites with collocated $\mathrm{SWE}$ and $z$ sensors (see Appendix B) only when both models and data yielded physical values ($0<\rho_{\mathrm{snow}}/\rho_{\mathrm{water}}<1$). Counts of “false non-snowpacks” (models show $z=0$ while $z>0$ in the data) and “false snowpacks” (models show $z>0$ while $z=0$ in the data) were recorded, along with instances of unphysical densities.
@@
Assessing physical consistency in the model directly per feature is nuanced due to strong correlations among the inputs, which confounds most interpretation methods like partial dependence, permutation, LIME, and tractable SHAP values (Molnar, 2022). To isolate feature effects on model output, we calculated first-order Accumulated Local Effects (ALE) plots (Apley and Zhu, 2020). This method shows the average change in model output accumulated over sequential bins of feature values. For feature $X$ at grid point $x_{i}$, data in the window $x_{i}-\Delta,x_{i}+\Delta]$ are evaluated setting $X=x_{i}-\Delta$ and $X=x_{i}+\Delta$, storing the average of the differences $\Delta M_{x_{i}}=\overline{M(x_{i}+\Delta)-M(x_{i}-\Delta)}$. The final (centered) ALE value $\bar{\Delta M}$ at $x_{i}$ is the sum over all $\Delta M_{x_{k}}$ with $x_{k}\leq x_{i}$, minus the average of all uncentered ALE values. This isolates changes solely from feature variations, and bins are defined by quantiles to ensure equal data instances in each window. The shape and slope of the ALE curve are more pertinent for interpreting physicality than the offset, especially when the feature distribution is skewed. The range of $\Delta\bar{M}$ over the feature indirectly measures feature importance in influencing model output, as the ALE value can be interpreted as a departure from the average model prediction. Further information for interpreting ALE plots can be found in Molnar (2022).
@@ 3.2 Density Timeseries
For validation and testing sites with collocated $z$ and SWE measurements (see Appendix B), Table 1 compares timeseries statistics for derived bulk density, and Fig. 7 displays estimates from the same data as Fig. 6. RMSE values reflect relative error since density is normalized by $\rho_{\mathrm{water}}$. Both SN17 and $\tilde{M}$ lack SWE discretization, inflating errors for small snowpacks compared to their assimilated counterparts. Neural unphysical densities source from a lack of constraint enforcing $\mathrm{SWE}=0\implies z=0$ (in both $\tilde{M},M$) or $z>\mathrm{SWE}$ (in $M$), though different constraint choices could eliminate this. Snow17 is roughly $3\times$ better at mitigating false snowpacks with lower RMSE (less extreme errors), but our models are roughly $3\times$ better at reducing false snow-absence and exhibit improved individual density predictions, crucial for estimating other snowpack properties. Thus, the advantage between models depends on which features are prioritized. Otherwise, the two perform similarly, despite $M$ being designed for snow depth while the comparable parameterization within Snow17 is explicitly formulated for density. Among evaluated sites, the best and worst performances by $M$ were better than those of SN17O.
@@ 3.4 Physical Behavior of Model
Figure 10: $M$ outputs for two snow states, with other inputs held constant. In each case one threshold (dz>zdz>-z or $dz/dt<0$ for $P_{\mathrm{snow}}=0$) is visible. “Shaded” or “masked” pixels (partially greyed out) indicate areas of the visible parameter space with ranges not represented in the data, either from being unphysical or unobserved.
@@ 3.6 Finer Resolution Predictions
Time units only appear as rates in $M$ via precipitation, the lower bound, and the output. By predicting rates $dz/dt$ instead of accumulated dzdz, $M$ can be evaluated at varying time-steps without retraining, by merely resetting the constraint function’s scaling of $1/\Delta t$ without altering the trained predictive weights (which were trained only once in section 3a, with $\Delta t$ = 1 day). This flexibility allows testing at resolutions up to data limits, which for Kühtai is a maximum resolution of 15-minute intervals. While sub-hourly time resolutions are rarer in larger earth models, they are used in land models and will become increasingly frequent as research advances in fine-scale land-atmosphere interactions (Schär et al., 2020; Ban et al., 2021). Similarly, while multiday resolutions are rare, data availability can necessitate their use. Benchmarking across scales above and below the daily resolution used for training offers insights into the model’s limits and potential applications. However, beyond a week, average input variables do not effectively capture critical input dynamics (e.g., monthly average $T_{\mathrm{air}}$ would not reliably indicate time above freezing), leading to suppressed variations that undermine meaningful outputs, and alternative model choices would be more appropriate. Results of repeating the timeseries generation with $M$ as described in Section 3f for the Kühtai data with different values of $\Delta t$ in the constraint but identical weights otherwise, spanning 15-minute to weekly resolutions, are shown in Fig. 12 and Table 2.
@@ Acknowledgements.
Then for a model output $p$ and any construction $f$ serving to threshold $p$, the bounds $\max(f,p)$ or $\min(f,p)$ can be explicitly implemented with a single depth-3 fixed-weight layer with no biases acting on input $p,f]^{\top}$ with ReLU activation, followed by an accumulation with no activation:
[…]
where ++ indicates $\max(f,p)$ and - indicates $\min(f,p)$, and the ReLU acts element-wise. Eq. 9 offers a simple formulation, but the symmetry of the max and min functions permits resonant structures—different weights yielding the same output. If the signs of $p$ or $f$ are unknown, both +f+f and f-f (or +p+p and p-p) must be passed through the ReLU, along with pfp-f, to preserve all necessary information. However, if $f$ is always nonnegative (or nonpositive), the layer can be simplified from three layers to two, as passing f-f (or +f+f) becomes redundant and its ReLU always evaluates to zero. Similarly, if the threshold obeys $f\geq C$ (or $p\geq C$) for some constant $C$, a similar reduction is possible by including a bias term along with $A_{2\pm}$ and $A_{1\pm}$. The same reductions apply if $p$ also exhibits similar properties. Fig. 13a illustrates the generalized structure for one-sided threshold constraints on a predictive component discussed in section 22.2.
@@
so the threshold can be explicitly implemented with a sequence of two fixed-weight layers containing no biases acting on input $p,f_{+},f_{-}]^{\top}$, followed by an accumulation with no activation:
[…]
@@
These constraint structures can adapt to any functional constraint $f$ of any input (including those independent of predictive inputs) or even the predictive component output pp. This versatile framework exhibits minimal computational overhead without increasing runtime complexity. This thresholds an output rather than imposing an invariance or zeroing of a derivative, but this constraint class is relevant to many physical or nonnatural systems. This provides a simple means for emulating many systems when paired with universally approximating predictive models, or integrating data-driven parameterizations into larger models while enforcing constraints like physics or conservation laws. Absolute boundaries also enhance stability in time-stepping by keeping outputs realistic.
@@
  • Only days with complete cases (no missing values in all variables) were extracted, and sequential differences $\Delta z_{i}=z_{i+1}-z_{i}$, $\Delta\mathrm{SWE}_{i}=\mathrm{SWE}_{i+1}-\mathrm{SWE_{i}}$, $p_{i}=AP_{i+1}-AP_{i}$ were calculated, where APAP is accumulated precipitation. Only values where $\Delta t_{i}=t_{i+1}-t_{i}=1$ day were kept, and the target $\frac{dz}{dt}_{i}\approx\frac{\Delta z_{i}}{\Delta t_{i}}$ created, as well as analogous $\frac{d\mathrm{SWE}}{dt}_{i}$ and $P_{i}=p_{i}/\Delta t_{i}$. Days with $P_{i}<0$ due to resetting of the water year were changed to $P_{i}$ = 0.
@@
  • Air temperatures in the continental US were bounded below by $40^{\circ}$C and air temperatures in Alaska were bounded below by $50^{\circ}$C instead of the instrument limit of $60^{\circ}$C to remove on average 1-2 individual suspect temperature spikes per site.